Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day11


Crew 103 Log Book for April 19, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: everybody is fine

Time Departed/Returned from EVA: 16h15-18h30

Brief Narrative of Field Mission Results:

 - This morning we tried to make our hot-air balloon fly but there was too much wind and we didn't succeed in getting it in the air.

 - The radiotelescope didn't detect traces of magnetic storms.

 - In the afternoon, we continued our experience about soil permeability in order to help Stacy Sklar from the Arizona State

University for her study.

 - We tried to dry our soil samples and found a humidity of about 1%

for the majority of the samples.

 - We're also trying to make an internet relay in order to allow our

max rover to follow us farther during our experience.

EVA Data/Interpretations:

We were wondering if the recent rains should modify the soil's

permeability, but the result was negative. The soil remains very dry.

Unfortunately the battery of our GPS turned off and we couldn't note

the coordinates of the three last samples.

Engineering/Hab Maintenance:

Nothing to mention

Plans for Tomorrow:

We hope there will be less wind so that our hot-air balloon can fly.

We will also change the LEDs from three blue LEDs to two blue and one red LED, so that we can study the influence of the red spectrum on their growth.

Inventory:

Support Requested:

Nothing for the moment


Commander's Report

Anne-Sophie Kalbfleisch Reporting

This morning we tried to send our beautiful hot-air balloon into air. Because of the high wind there were unwanted cold airflows and the warm air didn't succeed in sweeping our balloon into the air. So T. Windels and I did our exercises this morning: abdominal exercises, press-ups etc. These exercises did us a world of good!

T. Windels and S. Viroux tried to make an internet relay in order to allow our max rover follow us farther during our experiments. To do that we connected two Mac PC's together thanks to an ethernet cable.

We tried to configure the robot so that he could switch from one network to the other, in fact from the robot ADSL gateways to the network of our PC's. W. de Groot and I continued the measurements of our algae. We're reporting them in an excel file. At the end of our stay at MDRS we will interpret the results. We hope that we will have enough data because we received our experimental stuff very late from Fed-ex.

G. Van Roy, T. Windels and J. Offergeld did an EVA in order to continue our experience about the soil permeability and help Ms. Stacy Skylar (from Arizona State University) with her study. We were wondering if the recent rains would modify the soil's permeability, but the difference was almost inexpressible. We also dried our previous soil samples and found a humidity of about 1% for the majority of the samples.

See you tomorrow with new Martian news!


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,48

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,37

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 5 galons approximately

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 9,5

Hab Potable Water Tank Level (inches from bottom): 4

Potable Water Meter Reading: 45261,6

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: One of the two ventilators is on, I don't find the electric source for the other. The thermostat is at 50°F and the temperature inside is 90°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /

ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel consumed : 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 3 (Viking I):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 76

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 74


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well


General Engineering Notes/Comments: The battery of the ATV called Spirit seems to be down and the gearbox of the ATV called Opportunity has some problems...


EVA 7 Report

Jonas Offergeld Reporting

Time: 10.15 - 10.45 am

Crew Members: Offergled Jonas, Gregory Van Roy,Steve Viroux, Wouter de Groot

Site Location: In front of the MDRS

Transit Mode: By foot

Objectives: Test of hot air balloon

Results: there are too much wind so the air wasn't hot enough so the balloon doesn't take off.

We are going to test tomorrow if there are less wind.

Lessons Learned: The balloon doesn't fly when there are too much wind


EVA 8 Report

Gregory Van Roy Reporting

Time: 4.15 - 6.30 pm

Crew Members: Offergled Jonas, Windels Thibault, Gregory Van Roy

Site Location: Near the Hanksville road, we took GPS with us to determine coordinates but the batteries were flat after a few minutes so we don't have any coordinate for this EVA...

Transit Mode: With ATV

Objectives: Continue the permeability measures of the area

Results: We have 3 more data of permeability but without coordinate... So we will check coordinate tomorrow or Thursday and give you the results at this moment.

Lessons Learned: Check the GPS batteries before going in EVA!


Chef's Report

Gregory Van Roy Reporting

1a) Today was

A non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Nothing special today

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Beerbreak Beef Soup, Kung Fu Chicken Pasta, Mushrooms, Pesto sauce and pomegranate drink mix

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Kung Fu Chicken : We finished one big can, there still is one can left

Mushroom : 2/3 of the pack left

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

I read the mail of yesterday about the recipes but we don't really make

something new. We just follow the instructions for each product. When we made something complicated like bread, cake or something else, I always gave you the recipe.

5) Lessons learned (if any)

None

6) Comments/questions for Kim and Jean, the food study investigators

Tomorrow is the last cooking day of the mission because we will return "on Earth" Friday.